# Referring Expression Instance Retrieval and A Strong End-to-End Baseline

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 3875
@@ 1. Introduction
To address these limitations, we propose a new task: Referring Expression Instance Retrieval (REIR). REIR requires models to directly retrieve and localize the referred object instance from a gallery of images, based on a fine-grained natural language expression. To support this task, we introduce a new dataset, REIRCOCO, which provides high-quality instance-level expressions paired with precise bounding box annotations. As illustrated in Figure 1, existing datasets are not suitable for this task. TIR datasets like MSCOCO Captions (Chen et al., 2015) only contain image-level descriptions (e.g., A living room with a dog laying on a cushion) without object-level annotations, while REC datasets such as RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) rely on short, ambiguous expressions (e.g., Girl) that can refer to multiple targets in gallery settings. In contrast, REIRCOCO offers fine-grained and unambiguous expressions (e.g., The brown dog lies on a cushion near the fireplace with a red rug underneath), uniquely grounding specific object instances even in cluttered scenes. We construct REIRCOCO using an automated pipeline with two stages: generation and filtering. In the generation stage, we prompt a vision-language model to produce contextualized referring expressions for each object in the MSCOCO detection dataset (Lin et al., 2014b) and RefCOCO dataset. Each prompt incorporates the object’s attributes, spatial relations, and surrounding context to encourage relationally rich and instance-specific descriptions. In the filtering stage, a separate language model evaluates expression quality, removing samples that are ambiguous, inaccurate, or non-discriminative. The resulting dataset contains over 30,000 images and 200,000 uniquely annotated object instances, each paired with several fine-grained referring expressions that unambiguously correspond to a single target.
@@
To evaluate our approach, we construct strong baselines by combining and adapting representative TIR and REC methods in the REIR setting and these adapted methods perform poorly. We compare these baselines with our proposed CLARE under standardized settings and observe that CLARE achieves consistently superior performance on the REIR task. Extensive ablation studies further validate the effectiveness of both our proposed model and the REIRCOCO dataset. Furthermore, CLARE also performs competitively on traditional REC and TIR benchmarks, demonstrating strong generalization across vision-language tasks.
@@ 3.1. Task Definition
Formally, given a referring expression $x_{\text{text}}$ and a gallery of $N$ images $\{x_{\text{img}}^{i}\}_{i=1}^{N}$, the model is required to predict a bounding box $B=[x,y,w,h]$ for the referred object, which need to lie within the correct image $x_{\text{img}}^{i}$ that contains the target instance. The model receives no prior information about which image contains the object, it need to reason over all candidates to jointly determine both where and what to localize.
@@ 3.2. A New Dataset: REIRCOCO
As illustrated in Figure 2, REIRCOCO is built through a scalable two-stage pipeline powered by vision-language foundation models: generation and filtering. In the generation stage, we begin with object annotations from the COCO detection and RefCOCO datasets. For each target instance, we construct structured prompts that include its bounding box, category label, the image’s global caption, and contextual information about surrounding objects. These prompts are then passed to GPT-4o (Achiam et al., 2023) to generate 5–10 diverse referring expressions per instance. To reduce the risk of negative truths—i.e., cases where a query unintentionally matches multiple similar objects across different images—the prompts are designed to maximize referential uniqueness. In particular, the model is encouraged to incorporate relational cues and spatial context involving nearby entities (e.g., “the chair to the left of the woman in red”) so that the generated expressions uniquely describe the target instance within the broader visual scene. This design ensures that each query expression has only one valid match in the entire image gallery, aligning perfectly with the REIR task requirement. In the filtering stage, we use DeepSeek R1 (DeepSeek-AI et al., 2025) to verify the quality of the generated descriptions. Each image is paired with a verification prompt that includes the caption and candidate expressions. The model is instructed to reject expressions that are ambiguous or semantically inconsistent with the image context, and to retain those that are specific, grounded, and discriminative. This automated filtering ensures that the final dataset contains only high-quality, retrieval-friendly expressions. See Appendix for more details.
@@ 4.4. Training and Inference
Inference. CLARE supports scalable retrieval: instance-level features for all gallery images are precomputed. At test time, a query is encoded once and compared via dot product to all candidates, enabling efficient and accurate retrieval-localization without repeated fusion. This modular structure combines the flexibility of two-stream design with the fine-grained reasoning power of contrastive and expert-based modeling, making CLARE highly effective for the REIR task. See Appendix for more details.
@@ 5.2. Implementation Details
We use the ViT-B and ViT-L variants from SigLIP (Zhai et al., 2023) as our model backbone. For feature extraction, we used a Transformer encoder-decoder architecture following Deformable DETR (Zhu et al., 2020; Li et al., 2022b), consisting of 6 encoder layers and 6 decoder layers. The number of object queries $M$ is set to 900. The MORE module uses 1 shared expert for general semantics and a routed expert bank of 4 specialists. Each query activates 2 routed expert via a top-2 gating strategy, enabling targeted relation-aware representation learning.
@@ 5.3. Results on REIR
Qualitative Analysis. Figure 4 shows qualitative comparisons between our proposed CLARE model and a strong two-stage baseline composed of SigLIP and DeepSeek-VL2. The examples are sampled from the REIRCOCO test set and highlight challenging referring expressions that involve spatial relations, interactions, or multi-object reasoning(e.g., the brown dog lying on a cushion near the fireplace). The baseline often retrieves globally similar but incorrect images due to its lack of instance-level alignment. In contrast, CLARE accurately retrieves and localizes the target object by leveraging cross-image contrastive learning and relation-aware text encoding. These results qualitatively validate the effectiveness of our end-to-end instance-level alignment strategy, and highlight the advantages of modeling fine-grained semantics and inter-object relationships directly during retrieval.
@@ 5.5. Results on TIR
5.6. Ablation Study
Effect of Training Strategy. To assess the effect of our two-phase training pipeline, we compare four settings: (1) only pretrained, (2) directly finetuned on REIRCOCO without pretrain,, and (3) our staged setup, first pretrtined on object grounding datasets, then finetuned on REIRCOCO. As shown in table 4, both pretraining and REIRCOCO fine-tuning are essential. Pretraining provides grounding capabilities, while REIRCOCO tuning enables instance-level cross-image alignment. The full staged training achieves the best results, confirming the benefit of progressive supervision.
Effect of the MORE Module. We analyze the impact of different expert routing configurations in MORE. The number of shared experts is fixed to 1, while the number of routed experts and active selection varies. The table 5 shows that enabling routed experts leads to steady performance gains. Specifically, using 4 routed experts and selecting the top-2 per query gives the best result (56.85 BR@10), confirming that relation-aware routing enhances expression-level adaptation. In contrast, disabling routed experts (i.e., 0 selected) results in 53.28 BR@10. This demonstrates that modeling diverse relational cues, such as spatial positioning, co-occurrence patterns, or relative size, is essential for resolving fine-grained expression ambiguities. See Appendix for more details.
@@ A.1. Training Process
Stage 2: REIR Finetuning. We then finetune CLARE on the REIRCOCO dataset for 20 epochs. Each training batch samples 8 images with 16 object instances and their corresponding referring expressions. We retain the same multi-scale strategy as in pretraining, and continue using the StepLR scheduler. The model is trained with both retrieval loss ($\mathcal{L}_{\mathrm{retrieve}}$) and localization loss ($\mathcal{L}_{\mathrm{box}}$), with loss weights empirically set to 1.0 and 5.0 respectively. For efficient training on large-scale data, we follow Detic (Zhou et al., 2022) to adopt multi-task sampling across GPUs—each GPU processes different subsets of referring expressions per iteration.
Additional Techniques. To ensure stability and consistency across image domains, we apply standard data augmentations including random horizontal flipping, color jitter, and scale jittering. Mixed precision training is enabled to reduce memory footprint. During training, we also construct negative pairs by sampling mismatched expression-instance combinations to enhance contrastive discrimination. The total training takes approximately 100K iterations using 8 NVIDIA A100 GPUs with a batch size of 64.
@@ B.1. Origin Datasets
MS COCO Dataset The Microsoft Common Objects in Context (MS COCO) dataset (Lin et al., 2014b) is a large-scale image dataset developed by Microsoft to advance research in object recognition within complex, real-world scenes. It contains over 330,000 images, with more than 200,000 of them annotated in detail. These annotations span 80 object categories and 91 “stuff” categories, comprising over 1.5 million labeled instances. Each image contains an average of 7.7 object instances, making the dataset particularly suitable for tasks that require understanding of rich contextual information. MS COCO supports multiple computer vision tasks including object detection, instance segmentation, keypoint detection, and image captioning, and is widely used as a benchmark for evaluating model performance in real-world settings.
@@ C.1. CLARE Result
These observations highlight the challenge of grounding relational and compositional semantics in open-world settings, and suggest future work on enhancing reasoning over multi-object context and rare cases.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYsubmissionid: 3875
@@ 1. Introduction
To address these limitations, we propose a new task: Referring Expression Instance Retrieval (REIR). REIR requires models to directly retrieve and localize the referred object instance from a gallery of images, based on a fine-grained natural language expression. To support this task, we introduce a new dataset, REIRCOCO, which provides high-quality instance-level expressions paired with precise bounding box annotations. As illustrated in Figure 1, existing datasets are not suitable for this task. TIR datasets like MSCOCO Captions (Chen et al., 2015) only contain image-level descriptions (e.g., A living room with a dog lying on a cushion) without object-level annotations, while REC datasets such as RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) rely on short, ambiguous expressions (e.g., Girl) that can refer to multiple targets in gallery settings. In contrast, REIRCOCO offers fine-grained and unambiguous expressions (e.g., The brown dog lies on a cushion near the fireplace with a red rug underneath), uniquely grounding specific object instances even in cluttered scenes. We construct REIRCOCO using an automated pipeline with two stages: generation and filtering. In the generation stage, we prompt a vision-language model to produce contextualized referring expressions for each object in the MSCOCO detection dataset (Lin et al., 2014b) and RefCOCO dataset. Each prompt incorporates the object’s attributes, spatial relations, and surrounding context to encourage relationally rich and instance-specific descriptions. In the filtering stage, a separate language model evaluates expression quality, removing samples that are ambiguous, inaccurate, or non-discriminative. The resulting dataset contains over 30,000 images and 200,000 uniquely annotated object instances, each paired with several fine-grained referring expressions that unambiguously correspond to a single target.
@@
To evaluate our approach, we construct strong baselines by combining and adapting representative TIR and REC methods in the REIR setting, and these adapted methods perform poorly. We compare these baselines with our proposed CLARE under standardized settings and observe that CLARE achieves consistently superior performance on the REIR task. Extensive ablation studies further validate the effectiveness of both our proposed model and the REIRCOCO dataset. Furthermore, CLARE also performs competitively on traditional REC and TIR benchmarks, demonstrating strong generalization across vision-language tasks.
@@ 3.1. Task Definition
Formally, given a referring expression $x_{\text{text}}$ and a gallery of $N$ images $\{x_{\text{img}}^{i}\}_{i=1}^{N}$, the model is required to predict a bounding box $B=[x,y,w,h]$ for the referred object, which need to lie within the correct image $x_{\text{img}}^{i}$ that contains the target instance. The model receives no prior information about which image contains the object, it needs to reason over all candidates to jointly determine both where and what to localize.
@@ 3.2. A New Dataset: REIRCOCO
As illustrated in Figure 2, REIRCOCO is built through a scalable two-stage pipeline powered by vision-language foundation models: generation and filtering. In the generation stage, we begin with object annotations from the COCO detection and RefCOCO datasets. For each target instance, we construct structured prompts that include its bounding box, category label, the image’s global caption, and contextual information about surrounding objects. These prompts are then passed to GPT-4o (Achiam et al., 2023) to generate 5–10 diverse referring expressions per instance. To reduce the risk of negative truths—i.e., cases where a query unintentionally matches multiple similar objects across different images—the prompts are designed to maximize referential uniqueness. In particular, the model is encouraged to incorporate relational cues and spatial context involving nearby entities (e.g., “the chair to the left of the woman in red”) so that the generated expressions uniquely describe the target instance within the broader visual scene. This design ensures that each query expression has only one valid match in the entire image gallery, aligning perfectly with the REIR task requirement. In the filtering stage, we use DeepSeek R1 (DeepSeek-AI et al., 2025) to verify the quality of the generated descriptions. Each image is paired with a verification prompt that includes the caption and candidate expressions. The model is instructed to reject expressions that are ambiguous or semantically inconsistent with the image context, and to retain those that are specific, grounded, and discriminative. This automated filtering ensures that the final dataset contains only high-quality, retrieval-friendly expressions. See the Appendix for more details.
@@ 4.4. Training and Inference
Inference. CLARE supports scalable retrieval: instance-level features for all gallery images are precomputed. At test time, a query is encoded once and compared via dot product to all candidates, enabling efficient and accurate retrieval-localization without repeated fusion. This modular structure combines the flexibility of two-stream design with the fine-grained reasoning power of contrastive and expert-based modeling, making CLARE highly effective for the REIR task. See the Appendix for more details.
@@ 5.2. Implementation Details
We use the ViT-B and ViT-L variants from SigLIP (Zhai et al., 2023) as our model backbone. For feature extraction, we used a Transformer encoder-decoder architecture following Deformable DETR (Zhu et al., 2020; Li et al., 2022b), consisting of 6 encoder layers and 6 decoder layers. The number of object queries $M$ is set to 900. The MORE module uses 1 shared expert for general semantics and a routed expert bank of 4 specialists. Each query activates 2 routed experts via a top-2 gating strategy, enabling targeted relation-aware representation learning.
@@ 5.3. Results on REIR
Qualitative Analysis. Figure 4 shows qualitative comparisons between our proposed CLARE model and a strong two-stage baseline composed of SigLIP and DeepSeek-VL2. The examples are sampled from the REIRCOCO test set and highlight challenging referring expressions that involve spatial relations, interactions, or multi-object reasoning(e.g., the brown dog lying on a cushion near the fireplace). The baseline often retrieves globally similar but incorrect images due to its lack of instance-level alignment. In contrast, CLARE accurately retrieves and localizes the target object by leveraging cross-image contrastive learning and relation-aware text encoding. These results qualitatively validate the effectiveness of our end-to-end instance-level alignment strategy and highlight the advantages of modeling fine-grained semantics and inter-object relationships directly during retrieval.
@@ 5.5. Results on TIR
5.6. Ablation Study
Effect of Training Strategy. To assess the effect of our two-phase training pipeline, we compare four settings: (1) only pretrained, (2) directly finetuned on REIRCOCO without pretraining, and (3) our staged setup, first pretrained on object grounding datasets, then finetuned on REIRCOCO. As shown in table 4, both pretraining and REIRCOCO fine-tuning are essential. Pretraining provides grounding capabilities, while REIRCOCO tuning enables instance-level cross-image alignment. The full staged training achieves the best results, confirming the benefit of progressive supervision.
Effect of the MORE Module. We analyze the impact of different expert routing configurations in MORE. The number of shared experts is fixed to 1, while the number of routed experts and active selection varies. The table 5 shows that enabling routed experts leads to steady performance gains. Specifically, using 4 routed experts and selecting the top-2 per query gives the best result (56.85 BR@10), confirming that relation-aware routing enhances expression-level adaptation. In contrast, disabling routed experts (i.e., 0 selected) results in 53.28 BR@10. This demonstrates that modeling diverse relational cues, such as spatial positioning, co-occurrence patterns, or relative size, is essential for resolving fine-grained expression ambiguities. See the Appendix for more details.
@@ A.1. Training Process
Stage 2: REIR Finetuning. We then finetune CLARE on the REIRCOCO dataset for 20 epochs. Each training batch samples 8 images with 16 object instances and their corresponding referring expressions. We retain the same multi-scale strategy as in pretraining, and continue using the StepLR scheduler. The model is trained with both retrieval loss ($\mathcal{L}_{\mathrm{retrieve}}$) and localization loss ($\mathcal{L}_{\mathrm{box}}$), with loss weights empirically set to 1.0 and 5.0, respectively. For efficient training on large-scale data, we follow Detic (Zhou et al., 2022) to adopt multi-task sampling across GPUs—each GPU processes different subsets of referring expressions per iteration.
Additional Techniques. To ensure stability and consistency across image domains, we apply standard data augmentations including random horizontal flipping, color jitter, and scale jittering. Mixed precision training is enabled to reduce the memory footprint. During training, we also construct negative pairs by sampling mismatched expression-instance combinations to enhance contrastive discrimination. The total training takes approximately 100K iterations using 8 NVIDIA A100 GPUs with a batch size of 64.
@@ B.1. Origin Datasets
MS COCO Dataset The Microsoft Common Objects in Context (MS COCO) dataset (Lin et al., 2014b) is a large-scale image dataset developed by Microsoft to advance research in object recognition within complex, real-world scenes. It contains over 330,000 images, with more than 200,000 of them annotated in detail. These annotations span 80 object categories and 91 “stuff” categories, comprising over 1.5 million labeled instances. Each image contains an average of 7.7 object instances, making the dataset particularly suitable for tasks that require understanding of rich contextual information. MS COCO supports multiple computer vision tasks, including object detection, instance segmentation, keypoint detection, and image captioning, and is widely used as a benchmark for evaluating model performance in real-world settings.
@@ C.1. CLARE Result
These observations highlight the challenge of grounding relational and compositional semantics in open-world settings and suggest future work on enhancing reasoning over multi-object contexts and rare cases.